# Complex crystallographic reflection groups and Seiberg–Witten integrable systems: rank 1 case

### v2

@@ 3.3 Lax matrix
for a suitable matrices $L=L(p,q)$, $A=A(p,q)$. The Lax pair $L,A$ can be found following the method of [16]. We only need the Lax matrix $L$; it is calculated from the Dunkl operator (3.16) by adapting the recipe from [16]. Namely, let $\mathbb{C}(p,q)$ denote the space of (meromorphic) functions of $p,q\in\mathbb{C}$, with the $\mathbb{Z}_{m}$-action $s(p,q)=(\omega^{-1}p,\omega q)$. One considers $\mathbb{C}(p,q)\rtimes\mathbb{Z}_{m}$ acting on itself by left multiplication. If we use a vector-space isomorphism $\mathbb{C}(p,q)\rtimes\mathbb{Z}_{m}\simeq\mathbb{C}\mathbb{Z}_{m}\otimes\mathbb{C}(p,q)$, we can interpret the action of any element as a $\mathbb{Z}_{m}\times\mathbb{Z}_{m}$ matrix with entries from $\mathbb{C}(p,q)$. For example, multiplication by $q$ and $s$ are represented by the following matrices $Q$ and $S$, respectively:
@@
remain constant under the hamiltonian dynamics (3.1). Note that the hamiltonian $h(p,q)$ has degree $m$ in $p$. Since the coefficients $b_{i}=b_{i}(\alpha;p,q)$ have degree $<m$ in $p$ if $i<m$, we must have $b_{i}=b_{i}(\alpha)$. On the other hand,
@@ 6.1 Hitchin systems
Hitchin systems on algebraic curves are known as a rich source of complex integrable systems [27]. For curves of genus $g=0,1$ one needs to allow Higgs fields to have poles [28, 29, 30], and several many-body integrable systems have already been identified with Hitchin systems in that way. Our cases can be interpreted as Hitchin systems on the orbifold tori which are Riemann spheres with with three or four punctures as shown in figure 2. This agrees with the class-S theory description given in [31], and, equivalently, M-theory orbifold construction in [32].
@@
In the following, we explain how to obtain the Hitchin system description from the Lax presentation. First, by looking at the properties of the Lax matrix $L=L(\alpha)$, we can recognize $\tilde{\phi}:=L(\alpha)d\alpha$ as a $\mathbb{Z}_{m}$-equivariant Higgs field on $\mathcal{E}$. Namely, take ×$\mathbb{C}\times\mathbb{C}^{m}$ with the $\mathbb{Z}_{m}$-action $\omega.(\alpha,\xi)=(\omega^{-1}\alpha,S\xi)$, where $S$ is the matrix (3.18). This induces a $\mathbb{Z}_{m}$-action on the total space of $\tilde{E}:=\oplus_{i=0}^{m-1}\mathcal{L}_{\omega^{-i}q}$, where the line bundle $\mathcal{L}_{q}$ was defined in Remark 3.2. This gives a $\mathbb{Z}_{m}$-equivariant element $\tilde{\phi}\in\mathrm{End}\tilde{E}\otimes\Omega^{1}_{\mathcal{E}}(\sum x_{i})$. At the punctures $\tilde{\phi}$ has simple poles with $\mathrm{res}\,\tilde{\phi}|_{\alpha=x_{i}}$ diagonalisable, and with prescribed eigenvalues $\mu_{j}^{\vee}(x_{i})$. Further, by modifications at the fixed points one can make the pair $(\tilde{E},\tilde{\phi})$ $\mathbb{Z}_{m}$-invariant; the modified pair can then be obtained as a pullback of some Higgs bundle $(E,\phi)$ on $\mathbb{P}^{1}=\mathcal{E}/\mathbb{Z}_{m}$. That is how in general one relates equivariant Higgs bundles with Higgs bundles on orbifolds, which in turn are identified with (weakly) parabolic Higgs bundles [33] (see also [34, 35] for related studies). The conclusion is that the phase space of our integrable system through the map $(p,q)\mapsto L(\alpha)\mapsto\tilde{\phi}\mapsto\phi$ gets identified with an open subset of the moduli space $\mathcal{M}^{\vee}$ of parabolic $\mathrm{GL}_{m}$ Higgs bundles on $\mathbb{P}^{1}$. The parabolic data consists of the eigenvalues/eigenspaces of the residues of $\phi$ at the orbifolded points (punctures) in $\mathbb{P}^{1}$, with three punctures for $m=3,4,6$ and four punctures for $m=2$. (The definition of $\mathcal{M}^{\vee}$ also requires a choice of parabolic weights, but this is not important since we work on an open subspace of the moduli space.) The moduli space $\mathcal{M}^{\vee}$ carries a structure of a complex integrable system: the Hitchin fibration and Hitchin system. This identifies our integrable system with the Hitchin system on an open subset of $\mathcal{M}^{\vee}$. The Hitchin fibration is built from the family of spectral curves which coincide with our spectral curves $\Sigma_{z}^{\vee}=\widetilde{\Sigma}_{z}^{\vee}/\mathbb{Z}_{m}$. Since these have genus one, the fibers are isomorphic to $\Sigma_{z}^{\vee}$, $z\in\mathbb{C}$. The dynamics of the Hitchin system is linear along the fibers. Because the dynamics in $p,q$ coordinates along the elliptic curves $\Sigma_{z}$ is also linear, we conclude that
@@
Here is the sketch of how to observe (6.1) without resorting to Hitchin systems. Starting from the Lax matrix $L(p,q;\alpha)$ and its spectral curve $\widetilde{\Sigma}^{\vee}:\,\det(L-kI)=0$, we view the family of eigenlines $(L-kI)\ell=0$ parameterised by $(k,\alpha)\in\widetilde{\Sigma}$ as a line bundle $\mathcal{L}$ over $\widetilde{\Sigma}^{\vee}$. The dynamics $p(t),q(t)$ induces a dynamics $\mathcal{L}(t)$ on the Jacobian $\mathrm{Jac}(\widetilde{\Sigma}^{\vee})$. One then checks the following two properties: (1) the induced dynamics on $\mathrm{Jac}(\widetilde{\Sigma}^{\vee})$ is linear, and (2) $\mathcal{L}^{s}\cong\mathcal{L}$ for any $s\in\mathbb{Z}_{m}$. As a result, the linear motion along $\Sigma$ in the phase space is mapped onto a linear motion along a $\mathbb{Z}_{m}$-fixed subtorus in the Jacobian of $\widetilde{\Sigma}^{\vee}$, which is isomorphic to $\mathrm{Jac}(\widetilde{\Sigma}^{\vee}/\mathbb{Z}_{m})\cong\Sigma^{\vee}$. This implies the isomorphism (6.1).
@@ 6.2 Local systems, star-shaped quivers, and generalised DAHAs
On the Betti side, we have spaces of the monodromy data of the above Fuchsian systems. According to the general theory [54], these are modelled by multiplicative quiver varieties associated to the affine Dynkin quivers of type ${D}_{4}$, ${E}_{6,7,8}$. From yet another perspective, they appear in the work of Etingof, Oblomkov, and Rains [55] on generalised rank-one DAHAs. These varieties can be characterised as certain affine del Pezzo surfaces, see Sections 6 and 9 of [55].
@@ 6.3 Quantum curves and opers
As explained above, we can view the fibration $\{\Sigma^{\vee}_{z}\}$ on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$ as a Hitchin fibration over a punctured $\mathbb{P}^{1}$. Quantisation of $\Sigma^{\vee}_{z}$ gives a pencil of Fuchsian ODEs with prescribed local monodromy data. The monodromy around each puncture is semi-simple and has prescribed eigenvalues (with repetitions in cases $m=4,6$). If we write $G=\mathrm{GL}_{m}$ and denote by $M_{i}\in G$ the monodromy around $x=e_{i}$, then $M_{i}$ belongs to a particular semisimple conjugacy class $[\Lambda_{i}]:=\{g\Lambda_{i}g^{-1}\,|\,g\in G\}$. For example, for $m=6$ $\Lambda_{0}$ is a generic diagonal matrix while $\Lambda_{1,2}$ are of the form $\mathrm{diag}(a,a,b,b,c,c)$ and $\mathrm{diag}(d,d,d,e,e,e)$, respectively; the global monodromy in this case represents a point on the character variety
@@ Appendix A Elliptic functions and duality
$v_{l}(x,z)$ =$=\sum_{\{x_{i}\}}c_{l}(x_{i})e^{-\eta(\Omega_{l}x_{i})z}\varphi(x-x_{i},\Omega_{-l}z)\,,\qquad l\in\mathbb{Z}_{m}\setminus\{0\}\,,$ (A.4)
@@ A.2 Lie-theoretic interpretation of the duality
The duality transformation $c\mapsto c^{\vee}$ admits an interpretation in terms of the Weyl group action on quiver varieties. According to [67], any Fuchsian system gives rise to a representation of the deformed preprojective algebra $\Pi^{\lambda}$ of a star-shaped quiver $Q$. On the physics side, $\Pi^{\lambda}$ is defined in terms of the F-term equation for the corresponding mirror quiver theory, and the deformation parameter $\lambda$ can be identified with the Fayet-Iliopoulos parameters [68, 69, 70]. The number of legs of $Q$ equals the number of singular points of the system, and the leg lengths, deformation parameters $\lambda$, and the dimension vector are related to the local monodromy data at the singular points, see [67] for details. We are interested in the case of Fuchsian systems whose singularities and local monodromy are of the same type as for the quantum curves considered in Sec. 5, see Proposition A.21. Hence, they are systems of rank $m=2,3,4,6$ with regular singularities at the orbifold points of $\mathbb{P}^{1}=\mathcal{E}/\mathbb{Z}_{m}$. For each orbifold point $e_{r}$, let $\mathbb{Z}_{m_{r}}$ be the isotropy group of $e_{r}$. The local monodromy is required to be semisimple, with $m_{r}$ distinct eigenvalues of the same multiplicity $m/m_{r}$. Up to a twist, we may assume that at each singular point the product of local eigenvalues is $1$. Let $\ell$ denote the number of orbifold points, i.e., $\ell=4$ for $m=2$ and $\ell=3$ for $m=3,4,6$. The corresponding star-shaped quiver $Q$ will have the central node (indexed as 00) and $\ell$ legs, with nodes along the $r$th leg indexed as $[r,s]$, with $10\leq r\leq\ell-$ 1 and $11\leq s\leq m_{r}-$ 1. Hence, ignoring orientation of the edges, $Q$ is an affine Dynkin diagram of type $D_{4}$, $E_{6}$, $E_{7}$, ${E_{8}}$, see Fig. 4. Let $I$ be the set of the vertices of $Q$; they can be identified with simple roots $\alpha_{i}$ of the affine root system of type $Q$, so elements of $\mathbb{Z}^{I}$ are viewed as linear combinations of simple roots. The (affine) Weyl group $W$ of $Q$ acts on $\mathbb{Z}^{I}$; it is generated by simple reflection $s_{\alpha_{i}}$, $i\in I$. We set $\delta=\sum_{i\in I}n_{i}\alpha_{i}$ where the integers $n_{i}$ are indicated by the labels in Fig. 4, and note that $s_{\alpha_{i}}(\delta)=\delta$ for all $i\in I$.
@@ Appendix B Quantum Hamiltonians
[…]
@@
[…]
@@
[…]
@@
$\begin{split}\kappa_{1}&=(a_{2}\left(c_{2}-4b_{2}\right)+a_{4}+28b_{2}c_{2}+16b_{2}^{2}-6c_{2}^{2}\\ &\ \ +72b_{3}\hbar+\left(144b_{2}-42c_{2}\right)\hbar^{2})\wp\left(\omega_{1}\right)^{2}\end{split}$ $\begin{split}\kappa_{2}&=\omega^{2}\kappa_{1}\\ &\end{split}$ $\begin{split}\kappa_{3}&=\omega^{-2}\kappa_{1}\\ &\end{split}$
@@ C.2 Classical curves
The classical limit $\hbar=0$ of (C.2) can be written as $Q-zP=0$, where
@@
The pencil $Q-zP=0$ intersects the line $x-e_{i}w=0$ at two points $(e_{i}:\pm g_{i}\prod_{j\neq i}(e_{i}-e_{j}):1)$. To find its intersection with the line $w=0$, we set $x=1,w=0$ and get
@@
The classical limit $\hbar=0$ of (C.3) can be written as $Q-zP=0$, where
@@
Using the relation (2.48) (and setting $q=-1$), we see that this factorizes as
@@
Hence, the pencil $Q-zP=0$ passes through points $(1:-\mu_{j}(0):0)$. Therefore, we recognize this as a pencil of cubics from Sec. 4. Finally, the polynomial $Q$ can be rearranged as
@@
The classical limit of (C.5) is $Q-zP=0$, with
@@
The classical limit of (C.7) is $Q-zP=0$, with
@@
$Q_{5}$ =$=\{a_{3}b_{2}\left(e_{1}-e_{2}\right)w\left(x-e_{1}w\right)\left(x-e_{2}w\right)-a_{2}b_{3}\left(e_{1}-e_{2}\right)w\left(x-e_{1}w\right)\left(x-e_{2}w\right)$
@@
As explained above, each level set $h(p,q)=z$ of the hamiltonian is isomorphic, as an elliptic curve, to the dual curve $h^{\vee}(p,q)=z$, see (6.1). This will not be literally true for the curves as presented in this section. The reason is that our choice of $z$ as a parameter for elliptic pencils is not canonical, as it can be shifted by a constant (depending on the mass parameters). The same is true for the formulas for the hamiltonians $\hat{h}$ and $h$ in Sec. 2.6: they are defined up to an additive constant depending on the mass parameters. Due to this ambiguity, the isomorphism of curves in (6.1) becomes an isomorphism between a curve $Q-zP=0$ in the pencil and the dual curve curve $Q^{\vee}-z^{\vee}P=0$, where for $z^{\vee}=z+\mathrm{const}$ for a suitable constant depending on the mass parameters.
@@ Appendix D Algebraic integrability
This implies that the arithmetic progressions $\{j+\mu_{j}(x_{i})+m_{i}\mathbb{Z}_{\geq 0}\}$, $j=0,\dots,m_{i}-1$ do not overlap. Write $-n_{i}$ for the smallest number among $j+\mu_{j}(x_{i})$. Recall that $\mu_{j}(x_{i})$ sum to zero, therefore $n_{i}\geq 0$. Now consider the set
@@
with some $z\in\mathbb{C}$ determined by $t_{1},\dots,t_{n},\lambda$. The functions $\psi_{l}=\psi(\omega^{l}q)$ with $10\leq l\leq m-$ 1 span the solution space to the eigenvalue problem (D.8) for generic $z$.

### v3

@@ 3.3 Lax matrix
for suitable matrices $L=L(p,q)$, $A=A(p,q)$. The Lax pair $L,A$ can be found following the method of [16]. We only need the Lax matrix $L$; it is calculated from the Dunkl operator (3.16) by adapting the recipe from [16]. Namely, let $\mathbb{C}(p,q)$ denote the space of (meromorphic) functions of $p,q\in\mathbb{C}$, with the $\mathbb{Z}_{m}$-action $s(p,q)=(\omega^{-1}p,\omega q)$. One considers $\mathbb{C}(p,q)\rtimes\mathbb{Z}_{m}$ acting on itself by left multiplication. If we use a vector-space isomorphism $\mathbb{C}(p,q)\rtimes\mathbb{Z}_{m}\simeq\mathbb{C}\mathbb{Z}_{m}\otimes\mathbb{C}(p,q)$, we can interpret the action of any element as a $\mathbb{Z}_{m}\times\mathbb{Z}_{m}$ matrix with entries from $\mathbb{C}(p,q)$. For example, multiplication by $q$ and $s$ are represented by the following matrices $Q$ and $S$, respectively:
@@
remain constant under the hamiltonian dynamics (3.1). Note that the hamiltonian $h(p,q)$ has degree $m$ in $p$. Since the coefficients $b_{i}=b_{i}(\alpha;p,q)$ have degree <m<m in $p$ if $i<m$, we must have $b_{i}=b_{i}(\alpha)$. On the other hand,
@@ 6.1 Hitchin systems
Hitchin systems on algebraic curves are known as a rich source of complex integrable systems [27]. For curves of genus $g=0,1$ one needs to allow Higgs fields to have poles [28, 29, 30], and several many-body integrable systems have already been identified with Hitchin systems in that way. Our cases can be interpreted as Hitchin systems on the orbifold tori which are Riemann spheres with with three or four punctures as shown in Fig. 2. This agrees with the class-S theory description given in [31], and, equivalently, M-theory orbifold construction in [32].
@@
[…]
@@
[…]
@@ 6.2 Local systems, star-shaped quivers, and generalised DAHAs
On the Betti side, we have spaces of the monodromy data of the above Fuchsian systems. According to the general theory [54], these are modelled by multiplicative quiver varieties associated to the affine Dynkin quivers of type ${D}_{4}$, ${E}_{6,7,8}$. From yet another perspective, they appear in the work of Etingof, Oblomkov, and Rains [55] on generalised double affine Hecke algebras, or DAHAs. These varieties can be characterised as certain affine del Pezzo surfaces, see Sections 6 and 9 of [55].
@@ 6.3 Quantum curves and opers
As explained above, we can view the fibration $\{\Sigma^{\vee}_{z}\}$ on $T^{*}\mathcal{E}/\mathbb{Z}_{m}$ as a Hitchin fibration over a punctured $\mathbb{P}^{1}$. Quantisation of $\Sigma^{\vee}_{z}$ gives a pencil of Fuchsian ODEs with prescribed local monodromy data. The monodromy around each puncture is semi-simple and has prescribed eigenvalues (with repetitions in cases $m=4,6$). If we write $G=\mathrm{GL}_{m}$ and denote by $M_{i}\in G$ the monodromy around $x=e_{i}$, then $M_{i}$ belongs to a particular semisimple conjugacy class $\Lambda_{i}]:=\{g\Lambda_{i}g^{-1}\,|\,g\in G\}$. For example, for $m=6$ $\Lambda_{0}$ is a generic diagonal matrix while $\Lambda_{1,2}$ are of the form $\mathrm{diag}(a,a,b,b,c,c)$ and $\mathrm{diag}(d,d,d,e,e,e)$, respectively; the global monodromy in this case represents a point on the character variety
@@ Appendix A Elliptic functions and duality
$v_{l}(x,z)$ $=\sum_{\{x_{i}\}}c_{l}(x_{i})e^{-\eta(\Omega_{l}x_{i})z}\varphi(x-x_{i},\Omega_{-l}z)\,,\qquad l\in\mathbb{Z}_{m}\setminus\{0\}\,$, (A.4)
@@ A.2 Lie-theoretic interpretation of the duality
The duality transformation $c\mapsto c^{\vee}$ admits an interpretation in terms of the Weyl group action on quiver varieties. According to [67], any Fuchsian system gives rise to a representation of the deformed preprojective algebra $\Pi^{\lambda}$ of a star-shaped quiver $Q$. On the physics side, $\Pi^{\lambda}$ is defined in terms of the F-term equation for the corresponding mirror quiver theory, and the deformation parameter $\lambda$ can be identified with the Fayet-Iliopoulos parameters [68, 69, 70]. The number of legs of $Q$ equals the number of singular points of the system, and the leg lengths, deformation parameters $\lambda$, and the dimension vector are related to the local monodromy data at the singular points, see [67] for details. We are interested in the case of Fuchsian systems whose singularities and local monodromy are of the same type as for the quantum curves considered in Sec. 5, see Proposition A.21. Hence, they are systems of rank $m=2,3,4,6$ with regular singularities at the orbifold points of $\mathbb{P}^{1}=\mathcal{E}/\mathbb{Z}_{m}$. For each orbifold point $e_{r}$, let $\mathbb{Z}_{m_{r}}$ be the isotropy group of $e_{r}$. The local monodromy is required to be semisimple, with $m_{r}$ distinct eigenvalues of the same multiplicity $m/m_{r}$. Up to a twist, we may assume that at each singular point the product of local eigenvalues is $1$. Let $\ell$ denote the number of orbifold points, i.e., $\ell=4$ for $m=2$ and $\ell=3$ for $m=3,4,6$. The corresponding star-shaped quiver $Q$ will have the central node (indexed as 0) and $\ell$ legs, with nodes along the $r$th leg indexed as $[r,s]$, with $0\leq r\leq\ell-1$ and $1\leq s\leq m_{r}-1$. Hence, ignoring orientation of the edges, $Q$ is an affine Dynkin diagram of type $D_{4}$, $E_{6}$, $E_{7}$, $E_{8}$}, see Fig. 4. Let $I$ be the set of the vertices of $Q$; they can be identified with simple roots $\alpha_{i}$ of the affine root system of type $Q$, so elements of $\mathbb{Z}^{I}$ are viewed as linear combinations of simple roots. The (affine) Weyl group $W$ of $Q$ acts on $\mathbb{Z}^{I}$; it is generated by simple reflection $s_{\alpha_{i}}$, $i\in I$. We set $\delta=\sum_{i\in I}n_{i}\alpha_{i}$ where the integers $n_{i}$ are indicated by the labels in Fig. 4, and note that $s_{\alpha_{i}}(\delta)=\delta$ for all $i\in I$.
@@ Appendix B Quantum Hamiltonians
$\widehat{h}=\widehat{p}^{\,3}+(a_{2}\wp(q)+b_{2}\wp\left(q-\eta_{1}\right)+c_{2}\wp\left(q-\eta_{2}\right))\widehat{p}+\frac{1}{2}\left(a_{3}\wp^{\prime}(q)+b_{3}\wp^{\prime}\left(q-\eta_{1}\right)+c_{3}\wp^{\prime}\left(q-\eta_{2}\right)\right)\,.$ (B.2)
@@
In this case $\omega_{2}/\omega_{1}=\exp(\pi i/2)$, and we put $\omega_{0}=0$ for convenience. We have parameters $\mu_{j}(\omega_{i})$, $i=0,3$, and $\mu_{j}(\omega_{1})=\mu_{j}(\omega_{2})$, with $j=0,1,2,3$, and with $\mu_{j}(\omega_{1,2})=\mu_{j+2}(\omega_{1,2})$. Recall that $\sum_{j}\mu_{j}(x_{i})=0$ for each fixed point $x_{i}$. The hamiltonian has the form $\widehat{h}=\widehat{p}^{\,4}+A_{2}\,\widehat{p}^{\,2}+A_{3}\,\widehat{p}+A_{4}$ with the coefficients as follows (cf. [1, 37]):
@@
$\widehat{h}=\widehat{p}^{\,6}+A_{2}\,\widehat{p}^{\,4}+A_{3}\,\widehat{p}^{\,3}+A_{4}\,\widehat{p}^{\,2}+A_{5}\,\widehat{p}+A_{6}\,$, (B.9)
@@
[…]
@@ C.2 Classical curves
The classical limit $\hbar=0$ of (C.2) can be written as QzP=0Q-zP=0, where
@@
The pencil QzP=0Q-zP=0 intersects the line $x-e_{i}w=0$ at two points $(e_{i}:\pm g_{i}\prod_{j\neq i}(e_{i}-e_{j}):1)$. To find its intersection with the line $w=0$, we set $x=1,w=0$ and get
@@
The classical limit $\hbar=0$ of (C.3) can be written as QzP=0Q-zP=0, where
@@
Using the relation (2.48) (and setting q=1q=-1), we see that this factorizes as
@@
Hence, the pencil QzP=0Q-zP=0 passes through points $(1:-\mu_{j}(0):0)$. Therefore, we recognize this as a pencil of cubics from Sec. 4. Finally, the polynomial $Q$ can be rearranged as
@@
The classical limit of (C.5) is QzP=0Q-zP=0, with
@@
The classical limit of (C.7) is QzP=0Q-zP=0, with
@@
$Q_{5}$ $=\{a_{3}b_{2}\left(e_{1}-e_{2}\right)w\left(x-e_{1}w\right)\left(x-e_{2}w\right)-a_{2}b_{3}\left(e_{1}-e_{2}\right)w\left(x-e_{1}w\right)\left(x-e_{2}w\right)$
@@
As explained above, each level set $h(p,q)=z$ of the hamiltonian is isomorphic, as an elliptic curve, to the dual curve $h^{\vee}(p,q)=z$, see (6.1). This will not be literally true for the curves as presented in this section. The reason is that our choice of $z$ as a parameter for elliptic pencils is not canonical, as it can be shifted by a constant (depending on the mass parameters). The same is true for the formulas for the hamiltonians $\hat{h}$ and $h$ in Sec. 2.6: they are defined up to an additive constant depending on the mass parameters. Due to this ambiguity, the isomorphism of curves in (6.1) becomes an isomorphism between a curve QzP=0Q-zP=0 in the pencil and the dual curve curve $Q^{\vee}-z^{\vee}P=0$, where for $z^{\vee}=z+\mathrm{const}$ for a suitable constant depending on the mass parameters.
@@ Appendix D Algebraic integrability
This implies that the arithmetic progressions $\{j+\mu_{j}(x_{i})+m_{i}\mathbb{Z}_{\geq 0}\}$, $j=0,\dots,m_{i}-1$ do not overlap. Write $n_{i}$ for the smallest number among $j+\mu_{j}(x_{i})$. Recall that $\mu_{j}(x_{i})$ sum to zero, therefore $n_{i}\geq 0$. Now consider the set
@@
with some $z\in\mathbb{C}$ determined by $t_{1},\dots,t_{n},\lambda$. The functions $\psi_{l}=\psi(\omega^{l}q)$ with $0\leq l\leq m-1$ span the solution space to the eigenvalue problem (D.8) for generic $z$.